OpenStudy (anonymous):

Find two vectors v1 and v2 and whose sum is(3,0,2) , where v1 is parallel to(-4,-2,-4) while v2 is perpendicular to (-4,-2,-4)

OpenStudy (amistre64):

since v1 is parallel to .... then it is in fact equal to any scalar of .....

OpenStudy (amistre64):

and since v2 is perp to v1 by default; we need to pull a vector out of the plane that v1 is normaled to

OpenStudy (amistre64):

at any case: -4, -2 , -4 -4a, -2b, -4c ------------- -4-4a = 3 -2-2b = 0 -4-4d = 2 would be equations to fit in

OpenStudy (amistre64):

as well as 16a + 4b + 16c = 0 for the dot